OpenStudy (anonymous):

How do you simplify logarithms? I want to simplify log 8 = 4, what is the procedure? The answer is 2/3.

hartnn (hartnn):

can u please verify the question?1st u have not specified the base of logarithm,2nd there is nothing to simplify in log 8=4,unless u have base....

OpenStudy (anonymous):

Oops I mistyped my problem.. sorry Here it is :

OpenStudy (anonymous):

And I've never been taught logs before so an easy explanation would be fantastic, thanks :)

hartnn (hartnn):

http://openstudy.com/study?login#/updates/503a002fe4b0edee4f0d88db see the 12th property....if still in doubts,then ask again...

hartnn (hartnn):

u asked the question at right time,@TheViper just gave a tutorial on logarithms...

OpenStudy (anonymous):

use the property. \[\frac{ \log a }{ \log b } = \log(a - b)\]

OpenStudy (anonymous):

@Pratu043: Would that lead to the answer being 2/3 ? Sorry totally new at logarithms..

OpenStudy (anonymous):

so am i ....... TheViper is giving a tutorial, you can ask him.

OpenStudy (anonymous):

Thanks, will do :)

OpenStudy (anonymous):

Ahhhhh I still don't understand! What formula am I supposed to use?

hartnn (hartnn):

let me tell u how it leads to 2/3 i will be using his 6th property: \[\Large{6.)\log_ab^n=n\space \log_ab}\] and the fact that \[4=2^2 and8=2^3\] so \[\frac{\log 4}{\log 8}=\frac{\log 2^2}{\log 2^3}=\frac{2}{3}\frac{\log 2}{\log 2}=\frac{2}{3}\] got this??

OpenStudy (anonymous):

I understand the last bits just not the formula haha Thank you, this would have taken me eternity! :)